# MuSACo: Multimodal Subject-Specific Selection and Adaptation for Expression Recognition with Co-Training

### v1

@@ Abstract
Personalized expression recognition (ER) involves adapting a machine learning model to subject-specific data for improved recognition of expressions with considerable interpersonal variability. Subject-specific ER can benefit significantly from multi-source domain adaptation (MSDA) methods – where each domain corresponds to a specific subject – to improve model accuracy and robustness. Despite promising results, state-of-the-art MSDA approaches often overlook multimodal information or blend sources into a single domain, limiting subject diversity and failing to explicitly capture unique subject-specific characteristics. To address these limitations, we introduce MuSACo, a multimodal subject-specific selection and adaptation method for ER based on co-training. It leverages complementary information across multiple modalities and multiple source domains for subject-specific adaptation. This makes MuSACo particularly relevant for affective computing applications in digital health, such as patient-specific assessment for stress or pain, where subject-level nuances are crucial. MuSACo selects source subjects relevant to the target and generates pseudo-labels using the dominant modality for class-aware learning, in conjunction with a class-agnostic loss to learn from less confident target samples. Finally, source features from each modality are aligned, while only confident target features are combined. Our experimental results [^1] Our code is provided in suppl. materials and will be made public. on challenging multimodal ER datasets – BioVid and StressID – show that MuSACo can outperform UDA (blending) and state-of-the-art MSDA methods.
@@ 1 Introduction
Expression recognition (ER) has achieved significant success in computer vision due to its relevance in numerous real-world applications, such as pain estimation, stress monitoring, and affective computing [1, 2, 18, 37, 32]. These models can achieve a high level of performance under supervised settings or when provided with limited labeled data [25, 27]. However, their effectiveness mainly depends on the availability of annotated datasets, which are difficult to obtain, and often fail to generalize for subtle expressions and across diverse individuals due to variations in cultural, environmental, and individual expressiveness [13]. Unsupervised domain adaptation (UDA) [10, 20, 43] has emerged as a promising direction to address this challenge by adapting a model from a labeled source domain to an unlabeled target domain. In the context of ER, this facilitates personalization, where a model trained on labeled expression data from multiple source subjects is adapted to a new, unlabeled target subject. While existing UDA approaches [11, 18, 49] adapt to the unlabeled target dataset treated as a domain by either using single or blended labeled source subjects, recent advances in multi-source domain adaptation (MSDA) [45, 46] leverages information from multiple source subjects (domains). This encourages diversity in the source model for adaptation to an unlabeled target subject (domain). This method facilitates more precise and targeted adaptation strategies for developing a personalized ER model (see Fig. 1(a)).
@@
Despite these advances, the majority of subject-specific ER approaches are predominantly constrained to a single modality [11, 46]. However, human expressions and fine-grained affective cues are often difficult to capture reliably from a single modality alone [19, 47]. To address this limitation, we focus on a multimodal setting that leverages the complementary strengths of multimodal data. Multimodal ER (MER) has been widely studied in the literature [2, 1, 26, 34], typically outperforming unimodal systems by leveraging complementary information from multiple modalities. MER approaches integrate diversity from various modalities, such as visual, textual, audio, or physiological signals [7] to enhance model performance. In subject-based ER, most MER methods focus on fully or weakly labeled data [34, 26], hindering the ability of the model to have a better adaptation to unlabeled data. Nevertheless, there are some studies [15, 39] that introduced multiple modalities, specifically for EEG. Eye movement signals with a single source or combine multiple sources into a single domain (see Fig. 1(b)). Their reliance on a single source often fails to provide diversity in source domains. Furthermore, blending multiple sources into a single domain [16] cannot explicitly capture the unique subject-specific characteristics of each source subject, which is crucial for subject-specific target adaptation.
In this paper, we propose MuSACo, a novel approach for Multimodal Subject-Specific Selection and Adaptation for ER using Co-Training in an MSDA setting (Fig. 1(c)). Our method exploits the complementary information from multiple modalities while selecting the most relevant source subjects for the target. It takes advantage of the unique characteristics of each subject for better performance on the target subject (domain [^2] The terms domain and subject are used interchangeably.). There are two key challenges when considering multiple modalities for subject-specific MSDA: (1) effectively leveraging multiple source subject information within multimodal systems, and (2) achieving cross-domain and cross-modal alignment.
@@
To address the first challenge, we propose the selection of source subjects most relevant to the target subject. While prior work [45] has shown that selecting source subjects based on their relevance to the target improves adaptation, it relies solely on a single modality to guide the source selection process. In this paper, we propose a co-training-based strategy for selecting source subjects that leverage multiple modalities, as different modalities complement each other by providing unique characteristics from various perspectives. Thus, selecting sources using co-training benefits the adaptation process. Furthermore, to improve generalization across individuals, we encourage the model to focus on expression-specific features by disentangling identity-related information during source training.
@@
For the second challenge, we first focus on cross-domain alignment. State-of-the-art MSDA methods [17, 30] highlight the effectiveness of class-aware alignment in bridging dissimilar distributions. This typically involves generating pseudo-labels (PLs) for the target subject. In our proposed method, we extend this by leveraging co-training for the generation of target PLs. Specifically, PLs for unlabeled target data are generated by selecting predictions from the dominant modality (e.g., visual or physiological), chosen by their probability score and threshold. This ensures diversity in modality-specific feature representations and captures complementary aspects of the data. These PLs then guide class-aware alignment between source and target subjects, minimizing distribution mismatch. Although PLs allow improving the model performance, they heavily depend on the choice of a threshold to filter out unreliable samples. In particular, high threshold values ensure the sample quality. While setting a high threshold value often results in discarding potentially useful but less confident samples. To mitigate this, we introduce a class-agnostic loss that aligns non-confident target samples with the source. This encourages the unreliable samples to contribute to the learning process to improve the robustness of the model. For modality alignment, including a fusion module [26, 1] has proven to enhance MER performance. In MuSACo, a fusion module is added to concatenate features from different modalities for each selected source subject. For the target, only reliable samples (selected via co-training) are combined.
@@
The contributions of this paper are summarized as follows. (1) MuSACo: a novel multimodal MSDA method for accurate subject specific adaptation. It selects relevant source subjects from the target using co-training, leveraging complementary information across modalities. (2) An effective alignment process that relies on co-training to generate confident target PLs for class-aware loss, combined with class-agnostic loss to handle non-confident target samples. (3) Extensive experiments performed on challenging MER datasets (BioVid and StressID) shows the efficacy of MuSACo. It can outperform the UDA (blending) and state-of-the-art MSDA methods on this multimodal data.
@@ 2 Related Work
(a) Multi-Source Domain Adaptation: MSDA techniques for image classification can be classified into discrepancy-based methods [17, 23] that reduce the domain shift by estimating the discrepancy between source and target domains. [17], generate the target pseudo-labels by creating clusters for each class; the initialization will depend on the source domains. Contrastive learning [17, 30] methods focus on minimizing the disparity within the same class while maximizing with the opposite classes. Adversarial learning [48, 22], and self-supervised [36, 6] approaches, which train different classifiers for each source while generating pseudo-labels from the classifier agreement to adapt to the target domain. Recently, Zeeshan et al. [45] proposed a subject-based MSDA method for recognition expression that considers subjects as domains to tackle many sources (up to 77) for the adaptation process. Even though there are advancements in MSDA, they all depend on unimodal (visual) modality, which leads us to introduce multimodality in the MSDA setting for subject-based ER.
(b) Multimodal Expression Recognition: Multimodal ER is widely explored in the literature. Praveen et al. [26] proposed a cross-attention method that fused multiple modalities for expression recognition. Aslam et al. [1] presented an LUPI paradigm that learns from multiple modalities but uses only a single modality at the time of inference. Similarly, in [2], the author presented an optimal transport knowledge distilling method from different modalities. Multimodal methods also receive attention in domain adaptation for EEG and eye movement signals. Wang et al. [39] presented a variational autoencoder (VAE) multimodal method leveraging cycle consistency and adversarial learning losses to reduce the shift between domains and overcome the modality separation problem. Likewise, in [15], they adapt a model by reducing the discrepancy (CORAL loss [33]) between source and target distributions while adding an intra-inter domain loss for modality alignment for a cross-subject expression recognition using multimodal EEG signals. This paper exploits multiple modalities (visual and physiological) in MSDA for adapting to a new, unlabeled target subject.
@@
(c) Disentanglement in Expression Recognition: Disentanglement methods have been used in ER to suppress identity or pose information that interferes with ER. Jiang and Deng [14] proposed a GAN-based model with separate encoders for expression, identity, and pose, followed by a decoder and discriminator to guide expression reconstruction. Xie et al. [42] introduced TDGAN, which fuses identity from one image and expression from another. While effective for visual data, these GAN-based approaches rely on image reconstruction, making them unsuitable for non-visual modalities like physiological signals, where such reconstruction is not meaningful. Pichler et al. [24] proposed a KNIFE method that disentangled the information by estimating and minimizing the entropy between two estimators. Motivated by this strategy, our MuSACo adopts a similar formulation to decouple identity- and expression-specific representations.
@@
(d) Co-training for Domain Adaptation: Co-training is a well-established semi-supervised and unsupervised learning approach [3] that is used in many applications, such as semantic segmentation [41, 9], task decomposition [44], or in image classification [5]. It typically trains the model with modalities to produce more reliable pseudo-labels (PLs), while domain adaptation usually trains the model with a single modality that incorporates multiple classifiers to account for different views. Chen et al. [5] propose a method that introduces two different classifiers trained with different features controlled via a weighting mechanism while generating target pseudo-labels from the most confident classifier. MuSACo introduces a co-training method for multimodal data for the selection of source subjects. In addition, we select target confident PLs based on multiple classifiers trained on different modalities.
@@ 3.2 Training Backbones with Disentanglement
Two backbones are trained for each modality – visual $B_{\text{v}}$ and physiological $B_{\text{p}}$ – using multiple labeled source subjects $\mathcal{S}_{m}$. To learn expression-discriminative representations, each modality is supervised using a cross-entropy loss:
@@
Disentangled Identity-Information. Our approach is inspired by KNIFE [24], a differentiable entropy estimator, to disentangle identity information from expression-relevant features in each modality. We estimate the marginal entropy $H(\mathbf{h}_{m}^{\mathbf{s}})$ and the conditional entropy $H(\mathbf{h}_{m_{j}}^{\mathbf{s}}|\acute{Y}_{m_{j}}^{\mathbf{s}})$. For each sample in the source subjects, the identity labels $\acute{y}_{m_{j}}^{\textbf{s}}$ are converted into one-hot representations $\acute{Y}_{m_{j}}^{\textbf{s}}\in R^{k}$ and define the disentanglement loss as:
@@
The classification loss $\mathcal{L}_{m}^{s}$ and disentanglement loss $\mathcal{L}_{m}^{d}$ are jointly optimized to preserve task-relevant features while removing identity information [^3] More details are provided in Section 2 of the supplementary material..
@@ 3.3 Domain Adaptation via Co-training
(b) Generation of Target Pseudo-Labels (PLs). In MSDA, reliable PLs determine the performance of a model on an unseen target subject [45]. MuSACo benefits from the multimodal by selecting confident target PLs using co-training. Given two backbones, $B_{\text{v}}$ and $B_{\text{p}}$ We introduce samples $\mathbf{x}_{m}^{\textbf{t}}$ from $\mathbf{T}_{m}$ to produce softmax probabilities with respect to each modality, $p_{\text{v}}=\sigma(\mathbf{h}_{\text{v}}^{\textbf{s}})$ and $p_{\text{p}}=\sigma(\mathbf{h}_{\text{p}}^{\textbf{s}})$, taking the maximum probability $\widehat{p}=max(p_{\text{v}},p_{\text{p}})$ and applying a predetermined threshold $\tau_{pl}$ to select the confident samples, defined as:
@@
(c) Class-Aware Domain Alignment. We exploit class information to perform intra-class and inter-class alignment across domains. For intra-class alignment, samples from the same class are encouraged to reside in the same distribution space. Conversely, for inter-class alignment, samples from different classes are pushed further apart in the feature space. Class-aware alignment is computed using maximum mean discrepancy (MMD) [31], a statistical measure for estimating the disparity between two distributions. To sample out examples from the same class, we perform class-aware sampling on source and target subjects. The classes that belong to $\mathbf{T}^{l}_{m}$ were denoted as $\widetilde{C}$. We sample out examples from every subject belonging to multiple labeled sources. In each mini-batch, examples from every source and target are extracted for the same class. The intra-class discrepancy is defined as:
@@
(d) Class-Agnostic Domain Alignment. Most MSDA methods [17, 30] employ class-aware alignment to bridge dissimilar distributions, often relying on PLs for the target subject. However, these pseudo-labels are prone to noise, and while denoising techniques have been proposed [30, 43, 6, 28], they typically use pre-determined thresholds to retain only the most confident pseudo-labels (CPL). This reliance on thresholds can limit performance, especially in ER, where subtle expressions may lead to discarding useful but less confident samples. To address this, a domain-agnostic loss is introduced that aligns non-confident samples to ensure that all target samples contribute to the adaptation process. This approach maximizes the utilization of available data for improving the target adaptation. The domain agnostic loss is defined as:
@@
where $\mathbf{h}_{m_{a}}^{s}$ is the feature vector extracted from $\widetilde{\mathcal{S}}_{m}$ subjects that contain samples not belong to $\widetilde{C}$. The total alignment loss for both modalities is calculated as:
@@
(e) Modality Alignment. In multimodal settings, aligning modalities by adding a fusion network has proven to be very successful in enhancing model performance [26, 1]. In MuSACo method, to align the modalities, a fusion module $F$ is introduced that projects the combined modality embeddings into a shared feature space. To train $F$ for multiple source subjects, we use the embeddings $\mathbf{h}_{\text{v}}^{\textbf{s}}$ and $\mathbf{h}_{\text{p}}^{\textbf{s}}$ to perform feature concatenation $\mathbf{h}^{\textbf{s}}=\mathbf{h}_{\text{v}}^{\textbf{s}}\oplus\mathbf{h}_{\text{p}}^{\textbf{s}}$, and pass through $F$ to calculate a cross-entropy loss:

### v2

@@ Abstract
Personalized expression recognition (ER) involves adapting a machine learning model to subject-specific data for improved recognition of expressions with considerable interpersonal variability. Subject-specific ER can benefit significantly from multi-source domain adaptation (MSDA) methods – where each domain corresponds to a specific subject – to improve model accuracy and robustness. Despite promising results, state-of-the-art MSDA approaches often overlook multimodal information or blend sources into a single domain, limiting subject diversity and failing to explicitly capture unique subject-specific characteristics. To address these limitations, we introduce MuSACo, a multimodal subject-specific selection and adaptation method for ER based on co-training. It leverages complementary information across multiple modalities and multiple source domains for subject-specific adaptation. This makes MuSACo particularly relevant for affective computing applications in digital health, such as patient-specific assessment for stress or pain, where subject-level nuances are crucial. MuSACo selects source subjects relevant to the target and generates pseudo-labels using the dominant modality for class-aware learning, in conjunction with a class-agnostic loss to learn from less confident target samples. Finally, source features from each modality are aligned, while only confident target features are combined. Experimental results on challenging multimodal ER datasets – BioVid, StressID, and BAH – show that MuSACo outperforms UDA (blending) and state-of-the-art MSDA methods. Our code is available: https://github.com/osamazeeshan/MuSACo
@@ 1 Introduction
Expression recognition (ER) has achieved significant success in computer vision due to its relevance in numerous real-world applications, such as pain estimation, stress monitoring, and affective computing [aslam2023privileged, aslam2024distilling, li2018deep, waligora2024joint, sharafi2025disentangled]. These models can achieve a high level of performance under supervised settings or when provided with limited labeled data [praveen2020deep, rajasekhar2021deep]. However, their effectiveness mainly depends on the availability of annotated datasets, which are difficult to obtain, and often fail to generalize for subtle expressions and across diverse individuals due to variations in cultural, environmental, and individual expressiveness [jack2009cultural]. Unsupervised domain adaptation (UDA) [gretton2012kernel, long2016unsupervised, xu2018deep, guichemerre2024source] has emerged as a promising direction to address this challenge by adapting a model from a labeled source domain to an unlabeled target domain. In the context of ER, this facilitates personalization, where a model trained on labeled expression data from multiple source subjects is adapted to a new, unlabeled target subject. While existing UDA approaches [han2020personalized, li2018deep, zhu2016discriminative] adapt to the unlabeled target dataset treated as a domain by either using single or blended labeled source subjects, recent advances in multi-source domain adaptation (MSDA) [zeeshan2024subject, zeeshan2025progressive] leverages information from multiple source subjects (domains). This encourages diversity in the source model for adaptation to an unlabeled target subject (domain). This method facilitates more precise and targeted adaptation strategies for developing a personalized ER model (see Fig. 1(a)).
@@
Despite these advances, the majority of subject-specific ER approaches are predominantly constrained to a single modality [han2020personalized, zeeshan2025progressive]. However, human expressions and fine-grained affective cues are often difficult to capture reliably from a single modality alone [liu2021comparing, zhang2020emotion]. To address this limitation, we focus on a multimodal setting that leverages the complementary strengths of multimodal data. Multimodal ER (MER) has been widely studied in the literature [aslam2024distilling, aslam2023privileged, praveen2022joint, tzirakis2017end], typically outperforming unimodal systems by leveraging complementary information from multiple modalities. MER approaches integrate diversity from various modalities, such as visual, textual, audio, or physiological signals [dzedzickis2020human] to enhance model performance. In subject-based ER, most MER methods focus on fully or weakly labeled data [tzirakis2017end, praveen2022joint], hindering the ability of the model to have a better adaptation to unlabeled data. Nevertheless, there are some studies [jimenez2024cfda, wang2022multi] that introduced multiple modalities, specifically for EEG. Eye movement signals with a single source or combine multiple sources into a single domain (see Fig. 1(b)). Their reliance on a single source often fails to provide diversity in source domains. Furthermore, blending multiple sources into a single domain [jimenez2024mmda] cannot explicitly capture the unique subject-specific characteristics of each source subject, which is crucial for subject-specific target adaptation. In this paper, we propose MuSACo, a novel approach for Multimodal Subject-Specific Selection and Adaptation for ER using Co-Training in an MSDA setting (Fig. 1(c)). Our method exploits the complementary information from multiple modalities while selecting the most relevant source subjects for the target. It takes advantage of the unique characteristics of each subject for better performance on the target subject (domain [^1] The terms domain and subject are used interchangeably.). There are two key challenges when considering multiple modalities for subject-specific MSDA: (1) effectively leveraging multiple source subject information within multimodal systems, and (2) achieving cross-domain and cross-modal alignment.
@@
To address the first challenge, we propose the selection of source subjects most relevant to the target subject. While prior work [zeeshan2024subject] has shown that selecting source subjects based on their relevance to the target improves adaptation, it relies solely on a single modality to guide the source selection process. In this paper, we propose a co-training-based strategy for selecting source subjects that leverage multiple modalities, as different modalities complement each other by providing unique characteristics from various perspectives. Thus, selecting sources using co-training benefits the adaptation process. Furthermore, to improve generalization across individuals, we encourage the model to focus on expression-specific features by disentangling identity-related information during source training.
@@
For the second challenge, we first focus on cross-domain alignment. State-of-the-art MSDA methods [kang2020contrastive, scalbert2021multi] highlight the effectiveness of class-aware alignment in bridging dissimilar distributions. This typically involves generating pseudo-labels (PLs) for the target subject. In our proposed method, we extend this by leveraging co-training for the generation of target PLs. Specifically, PLs for unlabeled target data are generated by selecting predictions from the dominant modality (e.g., visual or physiological), chosen by their probability score and threshold. This ensures diversity in modality-specific feature representations and captures complementary aspects of the data. These PLs then guide class-aware alignment between source and target subjects, minimizing distribution mismatch. Although PLs allow improving the model performance, they heavily depend on the choice of a threshold to filter out unreliable samples. In particular, high threshold values ensure the sample quality. While setting a high threshold value often results in discarding potentially useful but less confident samples. To mitigate this, we introduce a class-agnostic loss that aligns non-confident target samples with the source. This encourages the unreliable samples to contribute to the learning process to improve the robustness of the model. For modality alignment, including a fusion module [praveen2022joint, aslam2023privileged] has proven to enhance MER performance. In MuSACo, a fusion module is added to concatenate features from different modalities for each selected source subject. For the target, only reliable samples (selected via co-training) are combined.
@@
The contributions of this paper are summarized as follows. (1) MuSACo: a novel multimodal MSDA method for accurate subject-specific adaptation. It selects relevant source subjects from the target using co-training, leveraging complementary information across modalities. (2) An effective alignment process that relies on co-training to generate confident target PLs for class-aware loss, combined with class-agnostic loss to handle non-confident target samples. (3) Extensive experiments performed on challenging MER datasets (BioVid, StressID, and BAH) show the efficacy of MuSACo. It can outperform the UDA (blending) and state-of-the-art MSDA methods on this multimodal data.
@@ 2 Related Work
(a) Multi-Source Domain Adaptation: MSDA techniques for image classification can be classified into discrepancy-based methods [kang2020contrastive, peng2019moment] that reduce the domain shift by estimating the discrepancy between domains. Contrastive-based approaches [kang2020contrastive, scalbert2021multi] generate target pseudo-labels using class-wise clustering, while adversarial and self-supervised methods [zhao2021madan, nguyen2021stem, venkat2020your, deng2022robust] train multiple source-specific classifiers and derive target pseudo-labels through classifier agreement for domain adaptation. Recently, zeeshan2024subject proposed a subject-based MSDA method that treats each subject as a domain to handle many sources. However, existing MSDA approaches remain unimodal (visual-only), motivating our introduction of multimodality for subject-based ER.
(b) Multimodal Expression Recognition: Multimodal ER has been widely studied. Cross-attention fusion has been proposed to combine multiple modalities for expression recognition [praveen2022joint], while LUPI-based methods [aslam2023privileged] and optimal transport distillation [aslam2024distilling] enable learning from multiple modalities but inference with fewer. In domain adaptation, multimodal approaches have been applied to EEG and eye-tracking, e.g., variational autoencoders with cycle consistency [wang2022multi] or CORAL-based alignment with intra/inter-domain losses [jimenez2024cfda]. Building on these, our work leverages visual and physiological modalities in a multi-source DA setting to adapt to unlabeled target subjects.
@@
(c) Disentanglement in Expression Recognition: Disentanglement methods suppress identity or pose factors that hinder ER. GAN-based models such as [jiang2022disentangling, xie2020facial] rely on image reconstruction, making them effective for visual data but unsuitable for non-visual modalities like physiology. In contrast, KNIFE [pichler2022differential] disentangles by minimizing entropy between estimators, a modality-agnostic approach. Motivated by this, MuSACo adopts a similar formulation to decouple identity from expression representations.
@@
(d) Co-training for Domain Adaptation: Co-training is a well-established semi-supervised and unsupervised learning approach [blum1998combining] that is used in many applications, such as semantic segmentation [xia2020uncertainty, gomez2023co], task decomposition [yang2021deep], or in image classification [chen2011co]. It typically trains the model with modalities to produce more reliable pseudo-labels (PLs), while domain adaptation usually trains the model with a single modality that incorporates multiple classifiers to account for different views. Chen et al. [chen2011co] propose a method that introduces two different classifiers trained with different features controlled via a weighting mechanism while generating target pseudo-labels from the most confident classifier. MuSACo introduces a co-training method for selecting source subjects from multimodal data. In addition, we select target confident PLs based on multiple classifiers trained on different modalities.
@@ 3.2 Training Backbones with Disentanglement
Two backbones are trained for each modality – visual $B_{\text{v}}$ and physiological $B_{\text{p}}$ – using multiple labeled source subjects $\mathcal{S}_{m}$. To learn expression-discriminative representations, each modality is supervised using a cross-entropy:
@@
Disentangled Identity-Information. Our approach is inspired by KNIFE [pichler2022differential], a differentiable entropy estimator, to disentangle identity information from expression-relevant features in each modality. We estimate the marginal entropy $H(\mathbf{h}_{m}^{\mathbf{s}})$ and the conditional entropy $H(\mathbf{h}_{m_{j}}^{\mathbf{s}}|\acute{Y}_{m_{j}}^{\mathbf{s}})$. For each sample in the source subjects, the identity labels $\acute{y}_{m_{j}}^{\textbf{s}}$ are converted into one-hot representations $\acute{Y}_{m_{j}}^{\textbf{s}}\in R^{k}$ and define the disentanglement loss as:
@@
The classification loss $\mathcal{L}_{m}^{s}$ and disentanglement loss $\mathcal{L}_{m}^{d}$ are jointly optimized to preserve task-relevant features while removing identity information [^2] More details are provided in Section 2 of the supplementary material.. Importantly, since $\mathcal{L}_{m}^{d}$ is entropy-based and modality-agnostic, it is applied independently to each modality (visual, physiological, and audio), ensuring that subject-specific identity cues do not interfere with expression-related representations and improving robustness across modalities.
@@ 3.3 Domain Adaptation via Co-training
(b) Generation of Target Pseudo-Labels (PLs). In MSDA, reliable PLs determine the performance of a model on an unseen target subject [zeeshan2024subject]. MuSACo benefits from the multimodal by selecting confident target PLs using co-training. Given two backbones, $B_{\text{v}}$ and $B_{\text{p}}$ We introduce samples $\mathbf{x}_{m}^{\textbf{t}}$ from $\mathbf{T}_{m}$ to produce softmax probabilities with respect to each modality, $p_{\text{v}}=\sigma(\mathbf{h}_{\text{v}}^{\textbf{s}})$ and $p_{\text{p}}=\sigma(\mathbf{h}_{\text{p}}^{\textbf{s}})$, taking the maximum probability $\widehat{p}=max(p_{\text{v}},p_{\text{p}})$ and applying a predetermined threshold $\tau_{pl}$ to select the confident samples, defined as:
@@
(c) Class-Aware Domain Alignment. We exploit class information to perform intra-class and inter-class alignment across domains. For intra-class alignment, samples from the same class are encouraged to reside in the same distribution space. Conversely, for inter-class alignment, samples from different classes are pushed further apart in the feature space. Class-aware alignment is computed using maximum mean discrepancy (MMD) [sejdinovic2013equivalence], a statistical measure for estimating the disparity between two distributions. To sample out examples from the same class, we perform class-aware sampling on source and target subjects. The classes that belong to $\mathbf{T}^{l}_{m}$ were denoted as $\widetilde{C}$. We sample out examples from every subject belonging to multiple labeled sources. In each mini-batch, examples from every source and target are extracted for the same class. The intra-class discrepancy is defined as:
@@
(d) Class-Agnostic Domain Alignment. Most MSDA methods [kang2020contrastive, scalbert2021multi] employ class-aware alignment to bridge dissimilar distributions, often relying on PLs for the target subject. However, these pseudo-labels are prone to noise, and while denoising techniques have been proposed [scalbert2021multi, xu2018deep, deng2022robust, ren2022multi], they typically use pre-determined thresholds to retain only the most confident pseudo-labels (CPL). This reliance on thresholds can limit performance, especially in ER, where subtle expressions may lead to discarding useful but less confident samples. To address this, a domain-agnostic loss is introduced that aligns non-confident samples to ensure that all target samples contribute to the adaptation process. This approach maximizes the utilization of available data for improving the target adaptation. The domain-agnostic loss is defined as:
@@
where $\mathbf{h}_{m_{a}}^{s}$ is the feature vector extracted from $\widetilde{\mathcal{S}}_{m}$ subjects that contain samples not belonging to $\widetilde{C}$. The total alignment loss for both modalities is calculated as:
@@
(e) Modality Alignment. In multimodal settings, aligning modalities by adding a fusion network has proven to be very successful in enhancing model performance [praveen2022joint, aslam2023privileged]. In MuSACo method, to align the modalities, a fusion module $F$ is introduced that projects the combined modality embeddings into a shared feature space. To train $F$ for multiple source subjects, we use the embeddings $\mathbf{h}_{\text{v}}^{\textbf{s}}$ and $\mathbf{h}_{\text{p}}^{\textbf{s}}$ to perform feature concatenation $\mathbf{h}^{\textbf{s}}=\mathbf{h}_{\text{v}}^{\textbf{s}}\oplus\mathbf{h}_{\text{p}}^{\textbf{s}}$, and pass through $F$ to calculate a cross-entropy loss: